OpenStudy (anonymous):

how do you find the derivative of y=e^(2/x)?

OpenStudy (amistre64):

Du e^u; u = 2/x

OpenStudy (anonymous):

I got y'= e^(2/x) * (-2/x^2). Is it the same thing as -2e^(2/x)?